A pentagon abcde in which sides , Ab and ed parallel each other and angle b c d 7:3:8 then find angle b c d?
Answer by ikleyn(53937)

.
Under the given condition,


< A + < E = 180°.  Hence, < B + < C + < d = 540° - 180° = 360°.


Also, < B = 7x,  < C = 3x  and < D = 8x,  where unknown "x" is the common measure of these 3 angles.


So we have  7x + 3x + 8x = 360°,  which implies  x = 360/18 = 20°.


Answer.  < B = 7x = 140°;  < C = 3x = 60°;  < D = 8x = 160°.

Solved.